# Hi-Gen: Generative Retrieval For Large-Scale Personalized E-commerce Search

### v1

@@ 3.2.2. Metric Learning Module
where is a positive hyper-parameter and is a distance function. In this paper, we use the L2-norm distance as the distance function. Meanwhile, the Adam optimization algorithm is adopted to solve Eq.6.
3.3. Document Identifiers (docIDs) Generation
@@
Using hierarchical clustering, the docID of an item can be acquired once the dense vector representation is given (Tay et al., 2022). However, this method yields an unintelligible docID that is devoid of real category information. To solve this problem, we designed a category-guided hierarchical clustering method for document identifier generation, and thus the generated docID has a strong and clear semantic meaning.
@@
3.4. Language Model With Position-aware Loss
@@
As shown in Eq.7, the position weight in the step (i.e.,) consists of the positional hierarchical, semantic, and efficient weights, which are represented as , , and , respectively. Greater weight is given to the earlier token by . causes a greater loss at the semantic layer for the irrelevant category. At the efficiency layer, attempts to generate tokens that are as efficient as the target token. , , and are hyperparameters that help to balance the effect of positional hierarchical, semantic, and efficient weight. The semantic layer’s length is , while the docID’s length is . The loss function for each token in the docID is represented by . is a distance function that judges how similar clusters are in efficiency, while determines whether two categories are semantically similar or not. represents the efficient score which is calculated in Algorithm 1. In our work, is defined as the difference of the Click-Through Rate (CTR), and bert-sentence-similarity333https://www.kaggle.com/code/eriknovak/pytorch-bert-sentence-similarity is utilized as . is a decay function, we set it to . In addition, we use the cross-entropy loss as .
3.5. Online Serving Optimization
@@
I2I retrieval (Linden et al., 2003a; Yang et al., 2020; Brin and Page, 1998; Qian et al., 2019; Linden et al., 2003b) uses the consumer’s recent behaviors as triggers to retrieve recommended items. Hi-Gen-I2I treats the limited items that are generated from Hi-Gen as triggers and uses an I2I retrieval method (e.g., Swing444Swing is an i2i algorithm proposed by Alibaba Group and widely used in the industry.) to perform the recall process. Thus, it also can expand recall results quickly.
@@
4. Experiments Setup
4.1. Datasets
We conduct experiments on two datasets (i.e.,AOL4PS (Guo et al., 2021) and AEDST555Our code and dataset will be publicly available.) to evaluate Hi-Gen for large-scale personalized search. Below, we describe the details of these two datasets.
AOL4PS is an open-source large-scale dataset tailored for personalized search research, that contains the user, the query, the corresponding clicked documents (URLs), and timestamps indicating the timeline of each user’s interaction history. Table 1 illustrates the respective quantities in both the training and testing sets, detailing the number of samples, users, queries, and documents in each. To meet the input requirements of Hi-Gen, we extract the first two tokens of each URL (e.g., www.spiritplay.org represented as www.spiritplay) as a category and construct a one-layer hierarchical category tree. Furthermore, we observed a 65 overlap between the training and testing sets in this dataset. To assess the generalization capability of the Hi-Gen model, we excluded samples from the testing set that had already appeared in the training set for zero-shot experimental analysis. The specific approach aligns with the methodology employed by PersonalTM (Lian et al., 2023).
AEDST is a billion-scale industrial dataset sourced from a real large-scale E-commerce search system, including the user, query, item, and click time information. Organized in chronological order of user behavior, we designate the samples from the first 7 days as the training set and the samples from the subsequent day as the testing set. The detailed distribution of quantities is shown in Table 1. A detailed analysis revealed that 87.33 of the samples in the testing set had never appeared in the training set. This phenomenon is attributed to the nature of large-scale E-commerce search systems, where users typically do not click or purchase the same item under the same query in a short period. Hence, the testing set of AEDST dataset can be directly utilized to evaluate the model’s generalization performance.
| Type | |||||
@@ 4.1. Datasets
4.2. Baselines
We chose the classical retrieval model BM25666https://github.com/dorianbrown/rank bm25 and the classical generative retrieval model DSI (Tay et al., 2022) as baselines. Additionally, for the AOL4PS dataset, we included the state-of-the-art model PersonalTM (Lian et al., 2023) as a baseline, which incorporates personalized information. For the AEDST dataset, we also compared different-sized T5 (Raffel et al., 2020) (Small777https://huggingface.co/t5-small and Base888https://huggingface.co/t5-base) models. Subsequent sections will provide a detailed analysis of the experimental results.
4.3. Implementation Details
All experiments are conducted on the Tesla A100 platform using the Adam optimizer. The representation learning model, metric learning model999https://github.com/KevinMusgrave/pytorch-metric-learning, and language model are assigned learning rates of 0.0001, 0.00001, and 0.00005, respectively. Correspondingly, their batch sizes are set at 512, 10, and 64. The representation and metric learning models yield embeddings with dimensions of 256 and 768, respectively. Fast-kmeans101010https://pypi.org/project/fast-pytorch-kmeans/ in which k is 10 is utilized to produce the docID. In cases where the cluster size exceeds 100, the hierarchical clustering algorithm is executed recursively; otherwise, termination occurs. The embedding of dimension 768 is harnessed to generate the docID for all DSI baselines, originating from a pre-trained BERT (Devlin et al., 2018)111111https://huggingface.co/bert-base-uncased model. For the proposed Hi-Gen, we utilized the following settings: (1) : 1.0; (2) : 0.1; (3) : 0.8; (4) : 0.1; (5) : 0.1; (6) : 1. Predictions are obtained by decoding the language model using the beam search methodology. Additionally, all experiments adopt k-fold cross-validation, with k set to 3.
4.4. Evaluation Metrics
@@
5. Experimental Results
5.1. Main Results
Table 2 illustrates the experimental outcomes for AOL4PS, providing compelling evidence that Hi-Gen surpasses other retrieval methods in large-scale personalized search systems. Hi-Gen beats the state-of-the-art model (PersonalTM) by 3.30 and 7.73 in Recall@1 and Recall@10, respectively. Especially, all generative retrievals including DSI outperform BM25 in Recall@1 and Recall@10.
@@ 5.1. Main Results
We compare Hi-Gen with DSI and BM25 in AEDST, as indicated in Table 3. Hi-Gen outperforms the best baseline model (BM25) on Recall@1, Recall@50, and Recall@100, improving by 4.62, 25.87, and 28.37, respectively. Compared to T5-Small, T5-Base has twice as many model transformer layers. Based on the research metrics of DSI and Hi-Gen, we could conclude that using T5-Base as the backbone outperforms T5-Small. It suggests that better results are obtained with more layers. Surprisingly, in AEDST, DSI performs even worse than BM25. The main reason is an abundance of zero-shot cases in AEDST. As seen in Table 5, this phenomenon also occurs in zero-shot circumstances of AOL4PS.
@@ 5.2. Ablation Study
In Table 4, the removal of any one of the four components of Hi-Gen results in a drop in performance, indicating the effectiveness of each component. Based on the results of experiments, we discover that category-guided hierarchical clustering is the most important part of Hi-Gen. When category-guided hierarchical clustering is eliminated, Recall@1 and Recall@10 drop by 21.22 and 22.71, respectively. Apart from category-guided hierarchical clustering, Hi-Gen also heavily relies on discriminative representation. Removing discriminative representation will result in a 17.00 and 18.30 decrease in Recall@1 and Recall@10, respectively. Furthermore, removing metric learning will decrease 8.02 and 8.86 in Recall@1 and Recall@10 in Hi-Gen. It proves that using discriminative representations learned through representation and metric learning, rather than BERT embedding, leads to significantly superior results for Hi-Gen. What’s more, Hi-Gen will reduce Recall@1 and Recall@10 by 6.51 and 7.16, respectively, when position-aware loss is not utilized.
@@
5.3. Zero-Shot Analysis
AEDST is a very sparse data set because zero-shot samples make up about 87.33 of the test data samples. The results of the experiment are shown in Table 3, where DSI performs far worse than Hi-Gen or even BM25. This result is in line with the findings published in (Lian et al., 2023). We conducted experiments on the AOL4PS zero-shot scenarios to better illustrate Hi-Gen’s effectiveness in such a scenario. Amazingly, Hi-Gen outperforms all baselines on any metric, as Table 5 illustrates. Moreover, compared to DSI, Hi-Gen yields notable gains in Recall@1 and Recall@10 of 36.78 and 41.07, respectively.
@@ 5.4. The Impact Of Hi-Gen Variants
5.5. Online Experiments
To examine Hi-Gen’s performance in a realistic E-commerce search system, we implemented it in a real-word large-scale online E-commerce search engine. To ensure accuracy, Hi-Gen-Cluster-2-I2I is utilized to produce a sufficient number of recall items. A live experiment is conducted for a week using an A/B testing framework. In the control group, 4.0 of users are chosen at random and shown search results that don’t come from Hi-Gen. An extra 4.0 users are chosen at random in the experiment group, and shown search results contain items produced by Hi-Gen. As indicated in Table 6, Hi-Gen increased RecallNum, CTR, CVR, PayCount, and GMV by 6.89, 0.83, 1.51, 0.92, and 1.42, respectively. It implies that Hi-Gen has significant advantages in terms of retrieving efficient items in a practical large-scale E-commerce search engine. Furthermore, Hi-GEN can meet online needs without increasing latency when compared to current complex online retrievals because it just generates a limited number of items by language model.

### v2

@@ III-B2 Metric Learning Module
where is a positive hyper-parameter and is a distance function. stands for the click label of the j-th sample in the PV, stands for the sample with the same click label as the j-th sample, and stands for the sample with different click label as the j-th sample. In this paper, we use the L2-norm distance as the function.
III-C Document Identifiers (docIDs) Generation
@@
Using hierarchical clustering, the docID of an item can be acquired once the dense vector representation is given [8]. However, this method yields an unintelligible docID that is devoid of real category information. To solve this problem, we designed a category-guided hierarchical clustering method for document identifier generation, and thus the generated docID has a strong and clear semantic meaning.
@@
III-D Language Model With Position-aware Loss
@@
As shown in Eq.7, the position weight in the step (i.e.,) consists of the positional hierarchical, semantic, and efficient weights, which are represented as , , and , respectively. Greater weight is given to the earlier token by . causes a greater loss at the semantic layer for the irrelevant category. At the efficiency layer, attempts to generate tokens that are as efficient as the target token. , , and are hyperparameters that help to balance the effect of positional hierarchical, semantic, and efficient weight. The semantic layer’s length is , while the docID’s length is . The loss function for each token in the docID is represented by . is a distance function that judges how similar clusters are in efficiency, while determines whether two categories are semantically similar or not. represents the efficient score which is calculated in Algorithm 1. In our work, is defined as the difference of the Click-Through Rate (CTR), and bert-sentence-similarity222https://www.kaggle.com/code/eriknovak/pytorch-bert-sentence-similarity is utilized as . is a decay function, we set it to . In addition, we use the cross-entropy loss as .
III-E Online Serving Optimization
@@
I2I retrieval [34, 35, 36, 37, 38] uses the consumer’s recent behaviors as triggers to retrieve recommended items. Hi-Gen-I2I treats the limited items that are generated from Hi-Gen as triggers and uses an I2I retrieval method (e.g., Swing333Swing is an i2i algorithm proposed by Alibaba Group and widely used in the industry.) to perform the recall process. Thus, it also can expand recall results quickly.
@@
IV Experiments Setup
IV-A Datasets
We conduct experiments on two datasets (i.e.,AOL4PS [39] and AEDST444Our code and dataset will be publicly available soon.) to evaluate Hi-Gen for large-scale personalized search. Below, we describe the details of these two datasets.
AOL4PS is an open-source large-scale dataset tailored for personalized search research, that contains the user, the query, the corresponding clicked documents (URLs), and timestamps indicating the timeline of each user’s interaction history. Table I illustrates the respective quantities in both the training and testing sets, detailing the number of samples, users, queries, and documents in each. To meet the input requirements of Hi-Gen, we extract the first two words of each URL (e.g., www.spiritplay.org represented as www.spiritplay) as a category and construct a one-layer hierarchical category tree. Furthermore, we observed a 65 overlap between the training and testing sets in this dataset. To assess the generalization capability of the Hi-Gen model, we excluded samples from the testing set that had already appeared in the training set for zero-shot experimental analysis. The specific approach aligns with the methodology employed by PersonalTM [14]. Since AOL4PS lacks relevance labels, we apply the relevance model, bert-sentence-similarity555https://www.kaggle.com/code/eriknovak/pytorch-bert-sentence-similarity, to evaluate the relationship between the query and documents in AOL4PS. The relevance label is 1 if the relevance score is more than 0.5, and 0 otherwise.
AEDST is a billion-scale industrial dataset sourced from AliExpress, a real large-scale E-commerce search system, including the user, the query, the corresponding items, the click and relevance label. Organized in chronological order of user behavior, we designate the samples from the first 7 days as the training set and the samples from the subsequent day as the testing set. The detailed distribution of quantities is shown in Table I. A detailed analysis revealed that 87.33 of the samples in the testing set had never appeared in the training set. This phenomenon is attributed to the nature of large-scale E-commerce search systems, where users typically do not click or purchase the same item under the same query in a short period. Hence, the testing set of AEDST dataset can be directly utilized to evaluate the model’s generalization performance.
| Type | Sample | Query | User | Doc | |
@@ IV-A Datasets
IV-B Baselines
To verify Hi-Gen’s effectiveness, we compare it with three existing retrieval methods: sparse retrieval, dense retrieval and generative retrieval on the AOL4PS dataset.
Sparse Retrieval. We chose BM25666https://github.com/dorianbrown/rank bm25 as the baseline of sparse retrieval model. BM25 is a widely-used sparse retrieval in information retrieval that retrieve documents based on their relevance to a query, factoring in term frequency and document length.
Dense Retrieval. Classic dense retrieval models, including P-Click[40], SLTB[41], HRNN[42], GRADP[43], are chosen to shows that Hi-Gen shows better performance than traditional retrieve-then-rank models. P-Click calculates the personalized score of a document as the percentage of clicks on this document from past queries that are identical to the input query. SLTB outputs personalized recalled items by utilizing diverse clicked-based or topic-based features. HRNN extends previous RNN-based session modeling with one additional GRU level that models the user activity across sessions and provides a seamless way of transferring the knowledge acquired on the long-term dynamics of the user interest to session-level. GRADP uses RNN-based networks to extract short-term and long-term user profiles from personalized historical information.
Generative Retrieval. In order to prove that Hi-Gen is an effective generative retrieval with better metrics, we compare it with the classic generative retrieval model DSI [8] and the state-of-the-art model PersonalTM [14]. DSI is a sequence-to-sequence generative retrieval model outperforms strong baselines by directly generating relevant document identifiers from queries without relying on an explicit index. It is a new information retrieval paradigm aims to address the limitations of dense retrieval framework based on the similarity score. PersonalTM proposes a novel decoder architecture that effectively leverages personalized context without increasing trainable parameters, and applys a hierarchical loss function to optimize the performance.
In the AEDST, we compare Hi-Gen with BM25, P-Click and DSI. We use T5-Small777https://huggingface.co/t5-small and T5-Base888https://huggingface.co/t5-base as the transformer backbones for DSI and Hi-Gen in order to examine the effects of various parameter sizes.
IV-C Implementation Details
All experiments are conducted on the Tesla A100 platform using the Adam optimizer. The representation learning model, metric learning model999https://github.com/KevinMusgrave/pytorch-metric-learning, and language model are assigned learning rates of 0.0001, 0.00001, and 0.00005, respectively. Correspondingly, their batch sizes are set at 512, 10, and 64. The representation and metric learning models yield embeddings with dimensions of 256 and 768, respectively. Fast-kmeans101010https://pypi.org/project/fast-pytorch-kmeans/ in which k is 10 is utilized to produce the docID. In cases where the cluster size exceeds 100, the hierarchical clustering algorithm is executed recursively; otherwise, termination occurs. The embedding of dimension 768 is harnessed to generate the docID for all DSI baselines, originating from a pre-trained BERT [44]111111https://huggingface.co/bert-base-uncased model. For the proposed Hi-Gen, we utilized the following settings: (1) : 1.0; (2) : 0.1; (3) : 0.8; (4) : 0.1; (5) : 0.1; (6) : 1. Predictions are obtained by decoding the language model using the beam search methodology. Additionally, all experiments adopt k-fold cross-validation, with k set to 3.
IV-D Evaluation Metrics
@@
V Experimental Results
V-A Main Results
Table II illustrates the experimental outcomes for AOL4PS, providing compelling evidence that Hi-Gen surpasses other retrieval methods in large-scale personalized search systems. Hi-Gen beats the state-of-the-art model (PersonalTM) by 3.30 and 7.73 in Recall@1 and Recall@10, respectively. Especially, all generative retrievals including DSI outperform baselines of sparse and dense retrievals in Recall@1 and Recall@10.
@@ V-A Main Results
In order to evaluate the effectiveness of Hi-Gen in a larger and more realistic dataset, we compare it with classic models of sparse, dense and generative retrievals: BM25, P-Click and DSI. As indicated in Table III, Hi-Gen outperforms the best baseline model (BM25) on Recall@1, Recall@50, and Recall@100, improving by 4.62, 25.87, and 28.37, respectively. Compared to T5-Small, T5-Base has twice as many model transformer layers. Based on the research metrics of DSI and Hi-Gen, we could conclude that using T5-Base as the backbone outperforms T5-Small. It suggests that better results are obtained with more layers. Surprisingly, in AEDST, P-Click and DSI perform even worse than BM25. The main reason is an abundance of zero-shot cases in AEDST. As seen in Table V, this phenomenon also occurs in zero-shot circumstances of AOL4PS.
@@ V-B Ablation Study
In Table IV, the removal of any one of the four components of Hi-Gen results in a drop in performance, indicating the effectiveness of each component. Based on the results of experiments, we discover that category-guided hierarchical clustering is the most important part of Hi-Gen. When category-guided hierarchical clustering is eliminated, Recall@1 and Recall@10 drop by 21.22 and 22.71, respectively. Apart from category-guided hierarchical clustering, Hi-Gen also heavily relies on discriminative representation. Removing discriminative representation will result in a 17.00 and 18.30 decrease in Recall@1 and Recall@10, respectively. Furthermore, removing metric learning will decrease 8.02 and 8.86 in Recall@1 and Recall@10 in Hi-Gen. It proves that using discriminative representations learned through representation and metric learning, rather than BERT embedding, leads to significantly superior results for Hi-Gen. What’s more, Hi-Gen will reduce Recall@1 and Recall@10 by 6.51 and 7.16, respectively, when position-aware loss is not utilized.
@@
V-C Zero-Shot Analysis
AEDST is a very sparse data set because zero-shot samples make up about 87.33 of the test data samples. The results of the experiment are shown in Table III, where DSI performs far worse than Hi-Gen or even BM25. This result is in line with the findings published in [14]. We conducted experiments on the AOL4PS zero-shot scenarios to better illustrate Hi-Gen’s effectiveness in such a scenario. Amazingly, Hi-Gen outperforms all baselines on any metric, as Table V illustrates. Moreover, compared to DSI, Hi-Gen yields notable gains in Recall@1 and Recall@10 of 36.78 and 41.07, respectively.
@@ V-D The Impact Of Hi-Gen Variants
V-E Online Experiments
To examine Hi-Gen’s performance in a realistic E-commerce search system, we implemented it in a real-word large-scale online E-commerce search engine. To ensure accuracy, Hi-Gen-Cluster-2-I2I is utilized to produce a sufficient number of recall items. A live experiment is conducted for a week using an A/B testing framework. In the control group, 4.0 of users are chosen at random and shown search results that don’t come from Hi-Gen. An extra 4.0 users are chosen at random in the experiment group, and shown search results contain items produced by Hi-Gen. As indicated in Table VI, Hi-Gen increased RecallNum, CTR, CVR, PayCount, and GMV by 6.89, 0.83, 1.51, 0.92, and 1.42, respectively. It implies that Hi-Gen has significant advantages in terms of retrieving efficient items in a practical large-scale E-commerce search engine. Furthermore, Hi-GEN can meet online needs without increasing latency when compared to current complex online retrievals because it just generates a limited number of items by the language model.